# Online Collision Risk Estimation via Monocular Depth-Aware Object Detectors and Fuzzy Inference

### v1

@@ III-B1 A risk-correlating metric as the output target
Technically, one can use any reasonable risk quantifier as the target. For instance, Wang et al. manually assigns discrete risk levels to driving scenes based on the number of nearby traffic agents [14]. Alternatively, Feth et al. calculates time-to-collision to other vehicles as a proxy [15]. In our work, we make use of a safety-oriented object detection metric that has been shown strongly correlated with actual AV collision rates in simulations [7].
@@
For an aggregated indication on an image frame, we similarly find the average of the USC scores for all ground-truth objects, resulting in $\widetilde{\mathsf{USC}}$. As USC is a scoring function for safety characterization (i.e., the higher, the better), we shall use the deficit from the full mark (i.e., $1-\widetilde{\mathsf{USC}}$) as the regression target for risk indication.
@@ III-B2 Constructing the FIS
Figure 3: Construction and output surfaces of the FIS via three different approaches, namely (a) knowledge-based crafting, (b) data-driven membership tuning, and (c) data-driven rule learning. For simplicity, the tilde above the variables for denoting frame-wise values are omitted.
Figure 4: Qualitative results of our monitor applied to the nuScenes dataset [11]. Blue boxes are from the 3D object detector [10] and yellow boxes are from the depth-based pipeline. Our monitor can successfully translate object detection alignment measures into collision risk indicator.
@@
Having formulated the targeted risk quantifier, we describe how a FIS can be constructed to map the alignment measures, $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$, to the target, $1-\widetilde{\mathsf{USC}}$. Notably, as the direction of depth discrepancy typically matters when estimating driving risks, we take out the absolute value operation from Eq. (3), allowing $\widetilde{\mathsf{RDD}}$ to range across $[-1,1]$.
@@
Now, to construct the FIS, we make use of MATLAB’s implementation of Mamdani system [30, 31] and formulate three different approaches. The first one is based on pure knowledge, specifying (i) high-level fuzzy rules and (ii) membership functions that map real values to linguistic terms for the input/output variables. The second one extends on the first one by having the membership functions optimized with a set of training data via local pattern search. Contrarily, the third one uses the training data to learn a rule set based on the crafted membership functions via global particle swarm optimization. More details about data preparation will be given in Sec. IV
@@
Fig. 3(a) provides the handcrafted rules based on different combination of the input variables as well as the membership functions (MFs) for $\widetilde{\mathsf{RDD}}$. To explain, low $\widetilde{\mathsf{RDD}}$’s mean that the predictions are much closer to the ego vehicle than the ground truths, and vice versa. Therefore, by intuition, we assign high risks to high $\widetilde{\mathsf{RDD}}$ cases. As for medium and low cases, we assign medium or low risks depending on the IoU values. Due to the space limit, MFs for $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{USC}}$ are not shown, but they take the same shapes on the range $[0,1]$. The resulting output surface of the FIS is displayed aside.
@@
Now, for tuning the membership function, we run the local pattern search algorithm for 1000 iterations using 5-fold cross-validation. It can be seen from Fig. 3(b) that the “high” and “medium”” functions are slightly squeezed to the right, and the “low” function is extended. Correspondingly, the output surface is more sharply decreasing at higher values of $\widetilde{\mathsf{RDD}}$ and has a larger plateau at the low end. To interpret, the decreasing part is likely because $1-\widetilde{\mathsf{USC}}$, based on its formulation, tends to tolerate errors and does not often give high values. On the other hand, the larger plateau possibly results from still certain risks when $\widetilde{\mathsf{RDD}}$ falls around $[-0.5,0.1]$ (i.e., the left end of the “medium” function before and after the tuning), as we have linked “low $\widetilde{\mathsf{RDD}}$ with medium risk” and “medium $\widetilde{\mathsf{RDD}}$ with low risk” (for medium $\widetilde{\mathsf{IoU}}$).
@@
The above two tendencies can be observed more apparently from the result of rule learning, for which we execute a global particle swarm optimization algorithm for 20 iterations, each having 100 evaluations. As a result, the optimizer directly tunes up the risk for two cases when $\widetilde{\mathsf{RDD}}$ is low and medium originally, while decreasing it for a high case. Consequently, the FIS has a smoother output surface than the handcrafted and membership tuning variants.
@@ IV Validation and Demonstration
In this section, we first conduct an open-loop validation with image sequences and then demonstrate our monitor in a closed-loop simulation.
@@ IV-A Object detector monitoring
TABLE II: Recalls of ground truths by $\tilde{\mathbf{S}}$ and $\tilde{\mathbf{P}}$ in various settings. ($\alpha$, $\beta$) denotes the matching thresholds for $\mathsf{IoU}$ and $\mathsf{RDD}$ respectively.
@@
We first do an evaluation comparing the objects retrieved from the depth-based pipeline against ground truths. As a benchmark, we take the ordinary 3D object detector. It can be seen in Tab. II that in nominal cases, the 3D object detector performs quite better than the depth-based pipeline, but the margin decreases once we allow for the more imprecise predictions. Notably, the depth-based pipeline appears to be more robust than the object detector in the case of sensor noises and snowy conditions, corresponding to recent findings showing the generalization and robustness potential of depth estimators [34].
Subsequently, Tab. III provides a detailed analysis on the depth-based object retrieval pipeline over the object detector’s predictions. For each scene condition, we first evaluate the object detector against ground truths and mark the true positives (TPs) as well as false negatives (FNs). Then, we run the depth-based object retrieval pipeline and check how many of the TPs and FNs are retrieved respectively. Corresponding to Tab. III, the object detector does not perform well in the noisy and snowy conditions, while the depth-based pipeline can recover almost half of the objects. On the contrary, the depth-based pipeline is weaker in night scenes, potentially due to the low illumination giving overall fewer hints for depth estimation.
Lastly, we examine qualitatively the risk estimating output. In Fig. 4, the top row contains a nominal driving scene, where in frame (i), there are multiple detection misses, but our monitor does not raise the risk level due to their distance. By contrast, in frame (ii), there are several safety-critical pedestrians not being detected, and hence the high risk. For the middle scene, a person in a wheelchair continues to be missed by the ordinary object detector. The depth-based pipeline can detect the person and trigger a high risk level. Finally, the bottom row shows a simulated snowy scene, where two jaywalkers are missed by the 3D object detector yet recovered by the depth-based pipeline.
@@ IV-B AV safeguarding
Finally, we demonstrate our monitor’s capability to safeguard an AV in unusual situations. To 1 briefly illustrate, a baseline AV has been designed for an automated valet parking (AVP) use case in the Prescan simulator [35]. In each mission, it has to drive from a hand-off zone to a parking slot while navigating through mixed traffic of other vehicles and pedestrians. In addition, to mimic real-world operating conditions, different effects such as low illumination, weather changes, and rare objects are added to the simulations. For instance, Fig. 5 shows a scenario with a cow that has not been seen by the baseline object detector, likely ending up as a false negative.
Now, to protect the AV against such challenging scenarios, we implement our perception monitor alongside a controller shield on top of the baseline AV stack. Whenever our monitor raises the risk level beyond 0.5, the controller shield starts to decelerate the AV. Fig. 5 depicts that our monitor successfully triggers the shield to decelerate the AV, which would have crashed into the cow. In general, we observe a successful collision avoidance rate beyond 80%. For a more lively demonstration with another scenario, readers can check a full use case video (where our run-time monitor is showcased from 9:16 minute) [36].
@@ V Conclusion
In this work, by observing and validating that the alignment measures between an ordinary object detector’s predictions and the retrieved objects from a depth-based pipeline can be correlated to the error measures of the object detector against ground truths, we implemented a monitoring framework to estimate ego vehicle’s collision risks based on the alignment measures. Specifically, we used a risk-correlating metric to optimize a fuzzy inference system for the run-time collision risk estimation. As a result, we are able to infer a relevant collision risk level using low-cost monocular camera images only. Experiments with the nuScenes dataset and a closed-loop simulator demonstrated the effectiveness of our approach.
Our work opens several interesting directions to explore. For instance, one may use additional signals such as scene flow estimates to find objects that move fast towards the ego vehicle. In addition, intuitive rules regarding object classification results can also be considered, e.g., a car object should normally occupy $x$ pixels at distance $y$. Lastly, how the monitoring results can be extended with active learning techniques to improve the perception results and hence the AV’s performance is under investigation.

### v2

@@ III-B1 A risk-correlating metric as the output target
Technically, one can use any reasonable risk quantifier as the target. For instance, Wang et al. manually assigns discrete risk levels to driving scenes based on the number of nearby traffic agents [14]. Alternatively, Feth et al. calculates time-to-collision to other vehicles as a proxy [15]. In our work, we make use of a safety-oriented object detection metric that has been shown to be strongly correlated with actual AV collision rates in simulations [7].
@@
For an aggregated evaluation on an image frame, we similarly find the average of the USC scores for all ground-truth objects, resulting in $\widetilde{\mathsf{USC}}$. As USC is a scoring function for safety characterization (i.e., the higher, the better), we use the deficit from the full mark, i.e., $1-\widetilde{\mathsf{USC}}$, as the regression target for risk indication.
@@ III-B2 Constructing the FIS
Figure 3: Construction and output surfaces of the FIS via three different approaches, namely (a) knowledge-based crafting, (b) data-driven membership tuning, and (c) data-driven rule learning. For simplicity, the tilde above the variables $\mathsf{RDD}$ and $\mathsf{IoU}$ for denoting frame-wise values are omitted.
@@
Having formulated the targeted risk quantifier, we describe how a FIS can be constructed to map the alignment measures, $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{RDD}}$, to the target, $1-\widetilde{\mathsf{USC}}$.
@@
Essentially, we make use of MATLAB’s Mamdani FIS toolbox [30, 31] and implement three different construction approaches. The first one is based on pure knowledge, specifying (i) high-level fuzzy rules and (ii) membership functions that map real values to linguistic terms for the input/output variables. The second one extends on the first one by having the membership functions optimized with a set of training data via local pattern search. Contrarily, the third one uses the training data to learn a rule set based on the crafted membership functions via global particle swarm optimization. More details about data preparation will be given in Sec. IV.
@@
Fig. 3(a) provides the membership functions (MFs) for $\widetilde{\mathsf{RDD}}$ and the handcrafted rules based on different combinations of the input variables. To explain, low $\widetilde{\mathsf{RDD}}$’s mean that the predictions are much closer to the ego vehicle than the ground truths, and vice versa. Therefore, by intuition, we assign high risks to high $\widetilde{\mathsf{RDD}}$ cases. As for medium and low cases, we assign medium or low risks depending on the IoU values. Due to the space limit, MFs for $\widetilde{\mathsf{IoU}}$ and $\widetilde{\mathsf{USC}}$ are not shown, but they take the same shapes on the range $[0,1]$. The resulting output surface of the FIS is displayed aside.
@@
Now, for tuning the MFs, we run the local pattern search algorithm for 1000 iterations using 5-fold cross-validation, as introduced in [32]. It can be seen from Fig. 3(b) that the “high” and “medium” MFs are slightly squeezed to the right, and the “low” MF is extended. Correspondingly, the output surface is more sharply decreasing at higher values of $\widetilde{\mathsf{RDD}}$ and has a larger plateau at the low-$\widetilde{\mathsf{RDD}}$ end. To interpret, the decreasing part is likely because $1-\widetilde{\mathsf{USC}}$, based on its formulation, tends to tolerate errors and does not often give high values. On the other hand, the larger plateau possibly results from still certain risks when $\widetilde{\mathsf{RDD}}$ falls around $[-0.5,0.1]$ (i.e., the left end of the “medium” function before and after the tuning), as we have linked “low $\widetilde{\mathsf{RDD}}$ with medium risk” and “medium $\widetilde{\mathsf{RDD}}$ with low risk” when $\widetilde{\mathsf{IoU}}$ is medium.
@@
The above two tendencies can be observed more apparently from the result of rule learning, for which we execute a global particle swarm optimization algorithm for 20 iterations, each having 100 evaluations [32]. As Fig. 3(c) shows, the optimizer directly tunes up the risk estimates for two cases when $\widetilde{\mathsf{RDD}}$ is low and medium. It also decreases the risk estimates in the high-$\widetilde{\mathsf{RDD}}$ high-$\widetilde{\mathsf{IoU}}$ case. To explain, when the object detector recognizes an object in a certain direction correctly (i.e., the high $\widetilde{\mathsf{IoU}}$), even though the depth estimate is far off (i.e., the high $\widetilde{\mathsf{RDD}}$), the AV would drive away from the direction of the object, resulting in smaller collision risks.
@@ IV Validation and Demonstration
Figure 4: Qualitative results of our monitor applied to the nuScenes dataset [11]. Blue boxes are from the 3D object detector [10], and yellow boxes are from the depth-based pipeline. Our monitor successfully translates object detection alignment measures into a useful collision risk indicator.
In this section, we first conduct an open-loop validation with image sequences and then demonstrate our monitor in a closed-loop simulation using the globally optimized variant from the above.
@@ IV-A Object detector monitoring
TABLE II: Recalls of ground truths $\tilde{\mathbf{G}}$ by depth-based objects $\tilde{\mathbf{S}}$ and object detector predictions $\tilde{\mathbf{P}}$ in various settings. ($\alpha$, $\beta$) denotes the matching thresholds for $\mathsf{IoU}$ and $\mathsf{RDD}$, respectively.
@@
We first evaluate the objects retrieved from the depth-based pipeline $\tilde{\mathbf{S}}$ against ground truths. As a benchmark, we take $\tilde{\mathbf{P}}$ from the ordinary 3D object detector. It can be seen in Tab. II that in nominal cases, the 3D object detector performs quite better than the depth-based pipeline, but the margin decreases once we allow for the more imprecise predictions. Notably, the depth-based pipeline appears to be more robust than the object detector in the case of sensor noises and snowy conditions, corresponding to recent findings showing the generalization and robustness potential of depth estimators [35].
Subsequently, Tab. III provides a detailed analysis on the depth-based object retrieval pipeline over the object detector’s predictions. For each scene condition, we first evaluate the object detector against ground truths and mark the true positives (TPs) as well as false negatives (FNs). Then, we run the depth-based object retrieval pipeline and check how many of the TPs and FNs are retrieved, respectively. Similar to Tab. II, the object detector does not perform well in noisy and snowy conditions, while the depth-based pipeline can recover almost half of the objects. On the contrary, the depth-based pipeline is weaker in night scenes, potentially due to the low illumination, which gives fewer hints for depth estimation.
Lastly, we examine qualitatively the risk estimating output. In Fig. 4, the top row contains a nominal driving scene, where in frame (i), there are multiple detection misses, but our monitor does not raise the risk level due to their distance. By contrast, in frame (ii), several safety-critical pedestrians are not detected, hence the high risk. For the middle scene, a person in a wheelchair continues to be missed by the ordinary object detector. The depth-based pipeline can detect the person and trigger a high-risk output. Finally, the bottom row shows a simulated snowy scene, where two jaywalkers are missed by the 3D object detector yet recovered by the depth-based pipeline.
@@ IV-B AV safeguarding
Finally, we demonstrate our monitor’s capability to safeguard an AV in unusual situations. To briefly illustrate, a baseline AV has been designed for an automated valet parking (AVP) use case in the Prescan simulator [36]. In each mission, the AV has to drive from a hand-off zone to a parking slot while navigating through mixed traffic of other vehicles and pedestrians. In addition, to mimic real-world operating conditions, different effects such as low illumination, weather changes, and rare objects are added to the simulations. For instance, Fig. 5 shows a scenario with a cow that has not been seen by the baseline object detector, likely ending up as a false negative.
Now, to protect the AV against such challenging scenarios, we implement our perception monitor alongside a controller shield on top of the baseline AV stack. Whenever our monitor raises the risk level beyond a threshold (empirically set as 0.5), the controller shield starts to decelerate the AV. Fig. 5 depicts that our monitor successfully triggers the shield to decelerate the AV, which would have crashed into the cow. In general, we observe a successful collision avoidance rate beyond 80%. For a more lively demonstration with another scenario, readers can check a full use case video (where our run-time monitor is showcased from 9:16 minute) [37].
@@ V Conclusion
In this work, by observing and validating that the alignment measures between an ordinary object detector’s predictions and the objects retrieved from a depth-based pipeline can be correlated to the errors of the object detector against ground truths, we implemented a monitoring framework to estimate ego vehicle’s collision risks based on the alignment measures. Specifically, we used a risk-correlating metric to optimize a fuzzy inference system for the run-time collision risk estimation. As a result, we are able to infer a relevant collision risk level using low-cost monocular camera images only. Experiments with the nuScenes dataset and a closed-loop simulator demonstrated the effectiveness of our approach.
Our work opens several interesting directions to explore. For instance, one may use additional signals such as scene flow estimates to find objects that move fast towards the ego vehicle. In addition, intuitive rules regarding object classification results can also be considered, e.g., a car object should normally occupy $x$ pixels at distance $y$. Lastly, we plan to investigate how the monitoring results can be extended with active learning techniques to improve the perception results and, hence, the AV’s performance.